Com Sci 221 — Programming Languages Assignment #2 — Winter 1995 Due Wednesday, 18 January

1. Do problem 1.1 c on p. 21 of the text.

2. Do problem 1.2 on pp. 21 of the text. Explain briefly but as precisely as possible the significance of the concepts Expression, Term, and Factor.

3. Do problems 2.2, 2.3, and 2.4 on p. 59 of the text, but only parts c, d, and e.